# Calibrating Generative Models

### v1

@@ D.2 Initializing the Base Diffusion Model to sample a Gaussian mixture
When the diffusion coefficient is chosen such that $\sigma(t)=\sqrt{\kappa(t)}$ and the linear coefficient $(\kappa(t))_{0\leq t\leq 1}$ satisfies $\kappa(t)\geq 0,\ \int_{0}^{1}\kappa(t)dt=+\infty$, then $\overset{\rightarrow}{p}(\boldsymbol{x}(0))\overset{d}{=}\mathcal{N}(\boldsymbol{0},\mathbb{I})$. We choose $\kappa(t)=t^{-1}$. Simply, (35) turns samples from $p_{\text{target}}$ into Gaussian noise. In practice, since the drift and diffusion coefficients defined by $\kappa(t)$ are unbounded (which violates the assumptions for existence and uniqueness of the solution to the SDE from Section˜D.1), we cap $\kappa(t)$ at some large $M$.
@@
Equation˜36 is useful since it tells how to generate samples from $p_{\text{target}}$: first draw samples from $\overset{\rightarrow}{p}(\boldsymbol{x}(0))\approx\mathcal{N}(\boldsymbol{x}(0)\ |\ \boldsymbol{0},\mathbb{I})$, then solve the SDE (36) numerically using Euler-Maruyama, for example. However, for general target distributions $p_{\text{target}}$, the score of the forward process is intractable, which yields the backward diffusion process (36) also intractable.
@@
D.3 Experimental Details
We perform all synthetic data experiments using Adam (Kingma & Ba, 2015) with default momentum hyperparameters $\beta=(0.9,0.999)$ and a cosine decay learning rate schedule (Loshchilov & Hutter, 2016). We perform $2\times 10^{3}$ CGM iterations for every experiment. We train on a single H100 GPU.
@@ D.3 Experimental Details
$u_{\theta}$ is a neural network with two hidden layers of dimension $256$ and SiLU activations, and $\log\overset{\rightarrow}{p}(\boldsymbol{x}(t))$ is the analytical score of the forward process that we described in Section˜D.2. In addition to $\boldsymbol{x}(t)$, we feed as input to $u_{\theta}$ a sinusoidal time embedding of dimension $32$. By initializing the weights of the output layer of $u_{\theta}$ to zero, we ensure that $p_{\theta}$ is initialized at $p_{\theta_{\text{base}}}$, the reversal of the forward diffusion process $\overset{\rightarrow}{p}$.
All synthetic data experiments are performed with batch size $M=10^{4}$. For CGM-relax, we select $\lambda$ by first performing calibration for each $\lambda$ on a log linear grid from $10^{0}$ to $10^{-3}$ with $10$ grid points. We choose the value of $\lambda$ for which $(\mathcal{L}^{\text{viol}})^{1/2}$ is reduced by a factor of $10$ and $\mathcal{L}^{\text{KL}}$ is the smallest. If no such value exists, we choose $\lambda$ for which $\mathcal{L}^{\text{KL}}$ is smallest. For CGM-reward, we compute $\widehat{\boldsymbol{\alpha}}_{N}$ using $N=10^{5}$ samples from $p_{\theta_{\text{base}}}$.
Appendix E Case Study Additional Details
In this section, we describe the experimental setup for our case studies with CGM-relax and CGM-reward from Section˜4. We provide explanations regarding the generative model classes $p_{\theta}$, CGM constraint functions $\boldsymbol{h}$ and targets $\boldsymbol{h}^{\ast}$, choice of CGM hyperparameters $\lambda$ and $N$, model architectures, and training procedures. We also include additional samples from our models before and after calibration.
Just as in our synthetic data experiments, we perform all experiments using Adam with default momentum hyperparameters $\beta=(0.9,0.999)$ and a cosine decay learning rate schedule. Additional common training details are shown in Table˜1. We train all models on a single H100 GPU.
@@ Appendix E Case Study Additional Details
Hyperparameter Genie2 ESM3-open TarFlow TinyStories-33M
@@
$\boldsymbol{x}$ Space $(\mathbb{R}^{100\times 3})^{100}$ $(\mathcal{S}^{100}_{4096})^{50}$ $\mathbb{R}^{256\times 256\times 3}$ $\mathcal{S}^{200}_{10{,}000}$
@@
Model parameters 1515M 1.4.4B 463463M 3333M
Training time (hrs) $48$ $2.3$ 33 $0.1$
E.1 Calibrating Genie2
@@
Since Genie2 is trained as the reversal of a discrete-time noising process (a DDPM, see Ho et al., 2020), we first convert the discrete-time denoising diffusion model to a (continuous-time) diffusion model. We achieve this by redefining the final timestep $T$ of the original denoising process to be time $1$ of the continuous-time process. To define the drift function, we take the DDPM transition mean defined at each time $t$ in the discrete-time process, divide it by $1/T=T$, and define the drift function to be equal to the resulting value in between times $t/T$ and $(t+1)/T$. The diffusion coefficient is similarly defined by the DDPM transition standard deviation at each time $t$ in the discrete-time process, but is instead scaled by $T^{1/2}$. This approach of converting the DDPM into a continuous-time diffusion model ensures that when the SDE is solved under the Euler-Maruyama scheme using a grid of $T$ timesteps (i.e. the original time grid used to define the DDPM), one samples from the original DDPM.
@@
For CGM-relax, we calibrate to k=99k{=}99 constraints on the bivariate CDF of alpha helix and beta strand content. And for CGM-reward, we calibrate to k=15k{=}15 constraints using $N=2.5\times 10^{4}$ samples from $p_{\theta_{\text{base}}}$. Since sampling from the Genie2 base model is time intensive, the sampling cost to compute $\widehat{\boldsymbol{\alpha}}_{N}$ (with small variance) is a downside to CGM-reward. Results reported in Figure˜3 are averages over $3$ trials, with two standard errors.
@@
Secondary structure annotation. As discussed in Section˜4.1, we measure the diversity of a collection of protein structures by computing the proportion of residues that lie in each of the three protein secondary structure types. For the CATH domains and Genie2, we perform annotations using the Biotite package (Kunzmann & Hamacher, 2018), which considers only $C_{\alpha}$ backbone atoms. For the CATH proteins (Sillitoe et al., 2021), we obtain the secondary structure distribution by annotating the domains collected and published by Ingraham et al. (2019).
@@ E.3 Calibrating TarFlow
Our motivation for choosing this problem was twofold. First, the quality of images generated by the base TarFlow model is high, such that a pre-trained classifier could attain high accuracy without fine-tuning. Second, we observe that the wildlife images generated by the base TarFlow model contained predominantly lions and leopards (Figure˜8), and rarely contained foxes or wolves. From $5\times 10^{3}$ samples annotated by GPT o5-mini, we computed animal proportions $\{0.3590,0.1260,0.0404,0.0256,0.2704,0.1752\}$.
@@
Although we require the model to state its confidence when labeling the images, we do not use these confidence scores for fine-tuning. The calibration function $\boldsymbol{h}(\boldsymbol{x})\in\mathbb{R}^{5}$ is a one-hot encoding of the first five classes. Since nearly all of the samples from the base TarFlow model are labeled as one of the six animals, we observe that choosing a six-dimensional constraint (i.e. adding cheetah) results in a poorly conditioned dual problem (7), since then the components of $\boldsymbol{h}$ are nearly linearly dependent. Our target is the uniform distribution over animals $\boldsymbol{h}^{\ast}=[.167,.167,.167,.167,.167]$.
@@
We perform calibration with regularization parameter $\lambda=10^{-4}$, and sample size $N=5\times 10^{3}$. We assess the success of calibration using two metrics: the total variation (TV) distance of the distribution over animal proportions (using $5\times 10^{3}$ annotated images) to the uniform distribution and the FID, computed using only the samples belonging to the wildlife class in the AFHQ training dataset. We use $5\times 10^{4}$ samples from the generative model to compute FID. It is important to note that the FID is an imperfect metric for assessing the quality of generated images since it will be lower for models whose animal class makeup is similar to that of the training distribution. To account for this, we evaluate CGM-relax on the maximum entropy reweighting of the training dataset to the uniform distribution over animal classes. In other words, we up or down weighted images belonging to a particular animal class in order to sample the six animals belonging to wildlife class with equal probability.
@@
Our best model, calibrated using CGM-relax with $\lambda=10^{-4}$, obtains class proportions $\{0.2248,0.0854,0.1750,0.1566,0.1668,0.1086\}$, evaluated using $5\times 10^{3}$ samples from the model; $0.0828$ of the samples were labeled as None. CGM-relax reduces the miscalibration error by nearly three times, from a TV distance of .306.306 to .108.108 (Figure˜9). However, the FID score increases from 15.9 to 21.9. CGM-reward is unsuccessful at calibrating the base model; both miscalibration and FID is roughly the same as the base model. Since CGM-reward remains close to the base model, we evaluate FID on the original AFHQ training dataset.
@@
In Figure˜8, we provide random generations from both the pre-trained and the model calibrated with CGM-relax with $\lambda=10^{-4}$. By examining samples from the calibrated model, we observe two axes along which sample quality worsens after calibration. First, some of the samples (those labeled as None) are dogs or cats, which lie outside the AFHQ wildlife class. Second, a greater proportion of samples depict blends of multiple animals.
@@
Figure 9: Calibrating TarFlow with CGM-relax reduces the TV distance of animal class labels to the uniform distribution by approximately three times. However, CGM-relax also produces fewer realistic samples, as measured by FID.
E.4 Calibrating TinyStories-33M
E.4.1 Autoregressive sampling and log-likelihoods
@@
Following the setup used by Eldan & Li (2023), we sample in the standard autoregressive fashion with no temperature scaling. To compute sequence log-likelihoods, we consider the prompt as given and ignore tokens generated after the first end-of-sequence (EOS) token. Let $m$ be the length of the prompt and $n$ the index of the first EOS token. Then sequence $\boldsymbol{x}$ has log-probability
@@
E.4.2 TinyStories constraint definition
To calibrate TinyStories-33M, we use a simple heuristic procedure to detect the gender of the story’s character associated with the profession in the prompt. The procedure returns $1$ for female, 1-1 for male, and 0 if the gender cannot be determined. Given a generated story, our procedure is as follows.
(1) Pronoun at sentence two. If the second sentence begins with a third-person singular pronoun, we assign gender based on that pronoun. This is common with our prompt templates, e.g., “Once upon a time there was a doctor named Sam. She was very kind…”.
@@ E.4.2 TinyStories constraint definition
E.4.3 TinyStories-33M figure details
Both panels of Figure˜5 were created using $20$ replicates per model with different Pytorch seeds, with points indicating the mean metric value across replicates. $2048$ stories were sampled per profession for each replicate, resulting in $14\times 2048=28672$ stories per model. Due to the high number of samples and replicates, two-times-standard-error-of-the-mean error bars are smaller than the markers, so are not shown.
@@ E.4.3 TinyStories-33M figure details
which takes maximum value $1$ if all samples used the same gender and minimum value 0 if there are an equal number of each gender. The overall miscalibration values shown on the y-axis of Figure˜5A were computed by taking the average miscalibration for the eight professions used for fine-tuning.
@@
This procedure was repeated 20 times for different sampling seeds to ensure variance from estimating $\boldsymbol{\alpha}$ had little effect on the outcome. The resulting error bars for the symmetrized KL are smaller than the marker size.
E.4.4 Example generations
@@
We provide example generations for four prompts before and after finetuning with CGM-relax with $\lambda=0.1$. To conserve space, more example generations are provided on GitHub at https://github.com/smithhenryd/cgm/tree/main/tinystories.
Listing 1: Samples from TinyStories-33M
Once upon a time there was a doctor named Jack. He was fit and strong, loved helping people.One day, Jack was working on an operation. He soon came across a little girl named Mary who was scared. Even though Jack told her not to worry, he said "hello" to her.Mary was still scared, and the doctor could tell her not to worry. He said that everything would be okay for her to do the operation. Mary felt a little bit better.Jack helped Mary and made sure she was fit. He even helped her little brother and made sure that she was always safe.At the end of the operation, Mary was smiling. She thanked Jack for his help and for helping her son. She had learned lessons about not being scared and having someone stronger than her.
Once upon a time there was a lawyer named Matt. He was very new and people stopped to talk to him. He came across a big building which said "Quiet down."Matt thought it was a bad idea but he was very curious. So he did as he was told and soon his house became quiet.All of a sudden he heard a voice and it said, "What are you doing?" It was the jail officer!Matt replied, "I thought it was a good idea to reverse the law in this house."The jail officer said, "Smart Jackie, you know a lot! This house is the law of the world."Matt smiled and said, "But it was fun to reverse the law."The jail officer nodded and said, "Yes it was. But the law hates them."Matt thanked the jail officer and went on his way.The moral of the story is that it pays to be creative, but also to be careful when you
Once upon a time there was a teacher named Miss Jane. She loved to answer any questions and her class was empty.One day, when Miss Jane was busy writing something on the top of her desk, she heard a voice coming from her kitchen floor. She looked down to see a small mouse squeezing a piece of cheese into the board.Miss Jane gasped and said, " mouse, that's naughty! Put that cheese down!"The mouse replied, "No, I won't. I'm making sure no one steals the cheese."Miss Jane was getting frustrated and said, "That's not right. You must ask before you litter into someone else's work. That's not nice."The mouse knew from the right place and quickly put the cheese back. Miss Jane saw what he had done and said, "That's the right thing to do!"Missy stared at him in amazement and said, "I suppose you won't tell everyone what to do with it again!"
Once upon a time there was a pilot named Bobby. He flew a plane on his travels.One day, Bobby was flying in his plane when he saw a tree fall. He crashed the plane right into the tree. Bobby felt so sorry. He called for help, but no one was around.Suddenly, he saw a 3-year-old child standing alone above the tree. Bobby flew to the child and offered to help. He told the little girl not to worry and helped him onto his plane. Together, they flew the plane back to its home in the tree. Bobby and the 3-year-old child became good friends after that until finally Bobby had to fly on time.The end.
@@
Once upon a time there was a doctor named Susie. She was a very patient and kind lady who always supported others. Every day she took her to see lots of patients. She always tried to help them feel better. She was always so kind and always willing to lend a hand.One day, a patient called Sally came to see Susie. Sally was very sick and needed help. Susie was always so gentle and caring, not taking extra time to help others. When Sally asked her if she was okay, Susie just smiled and reassured her that she would be alright soon. She was so patient with Sally as she played and chased around the toy store.It was so nice to see how patient and caring Susie is between helping others. She continues to be sure to always lend a hand whenever she can and remember to be there for her when feeling ill.
Once upon a time there was a lawyer named Cass. Cass and Tim went to a big city. They decided to sit on the corner. It was a little dizzy when they sat because they were too dizzy.At the office, Mommy and Daddy talked to the boss. They asked the lawyer for help. They listened to his ideas and his ideas. The lawyer gave them big smiles. He showed them how to mix some ingredients together to make a cookie. The lawyer's ideas and his ideas fit the cookie dough. His ideas were so neat and the cookie dough was extra yummy! The lawyer thanked him and said, "Thanks for your help. You kept your ideas nice and happy." Cass and Tim smiled as they watched the lawyer leave the office. They waved goodbye as they walked away, still feeling happy each
Once upon a time there was a teacher named Mr. Jam and Mrs. Bunny. Mr. Bunny has very cool haircuts that everyone wanted. One day Mrs. Bunny said, "Tomorrow we will have a new surprise."The next morning, Mrs. Bunny came over to Rob's house. When they got to the house Mrs. Bunny said, "I have a new surprise for you!" She pulled out a big bicycle ticket.Mr. Bunny said, "This is for you my grandkids. Now settle down and take it for a ride." The little bunnies were so happy!They sat quietly and quietly, taking turns on the bicycle. Then Mrs. Bunny gave them a surprise, a cool drink. The little bunnies were so excited. They each had a cool drink!All the bunnies were so happy to be outside on such a cool day. And they thanked Mrs. Bunny for the special surprise.
Once upon a time there was a pilot named Judy. She loved to measure everything: trees, houses, farms, and anything else. She was really careful to measure each inch so that every time that she got a bit closer to the number. One day Judy was flying to measure a star in the sky. She drew a line through the line with her finger and measured it for a long time. She was almost done when something terrible happened. One of the letters flew too close and hit Judy in the face. It hurt a lot!Judy was very frightened and ran away from the window to escape from the bad hit. But unfortunately the bad letter kept rolling closer and closer until it was right at the edge of the world. Judy was so scared and upset she couldn't believe it! She had been measured and measured, but still got past the bad letter.The bad letter that Judy had measured was gone forever and Judy was left feeling very sad and lonely.

### v2

@@ D.3 Exact Sampling from a Gaussian Mixture Model
When the diffusion coefficient is chosen such that $\sigma(t)=\sqrt{\kappa(t)}$ and the linear coefficient $(\kappa(t))_{0\leq t\leq 1}$ satisfies $\kappa(t)\geq 0,\ \int_{0}^{1}\kappa(t)dt=+\infty$, then $\overset{\rightarrow}{p}(\boldsymbol{x}(0))\overset{d}{=}\mathcal{N}(\boldsymbol{0},\mathbb{I})$. Such a noise schedule satisfies the independence assumption in Section˜D.2. We choose $\kappa(t)=t^{-1}$. Simply, (34) turns samples from $p_{\text{target}}$ into Gaussian noise. In practice, since the drift and diffusion coefficients defined by $\kappa(t)$ are unbounded (which violates the assumptions for existence and uniqueness of the solution to the SDE from Section˜D.1), we cap $\kappa(t)$ at some large $M$.
@@
Equation˜35 is useful since it tells how to generate samples from $p_{\text{target}}$: first draw samples from $\overset{\rightarrow}{p}(\boldsymbol{x}(0))\approx\mathcal{N}(\boldsymbol{x}(0)\ |\ \boldsymbol{0},\mathbb{I})$, then solve the SDE (35) numerically using Euler-Maruyama, for example. However, for general target distributions $p_{\text{target}}$, the score of the forward process is intractable, which yields the backward diffusion process (35) also intractable.
@@
Appendix E Additional Experimental Details
In this section, we provide further details regarding our experiments with CGM-relax and CGM-reward, both on synthetic data (Section˜3) and in the real-data case studies (Section˜4). We provide explanations regarding the generative model classes $p_{\theta}$, CGM constraint functions $\boldsymbol{h}$ and targets $\boldsymbol{h}^{\ast}$, choice of CGM hyperparameters $\lambda$ and $N$, model architectures, and training procedures. We also include additional samples from our models before and after calibration.
We perform all experiments using Adam with default momentum hyperparameters $\beta=(0.9,0.999)$ and a cosine decay learning rate schedule. Additional common training details are shown in Table˜1. We train all models on a single H100 GPU.
E.1 Synthetic Data Experiments
For all synthetic data experiments we perform $2\times 10^{3}$ iterations of CGM-relax or CGM-reward. All results reported in Figure˜3 are averages over $10$ iterates, with two standard errors.
@@ E.1 Synthetic Data Experiments
$u_{\theta}$ is a neural network with two hidden layers of dimension $256$ and SiLU activations. $\log\overset{\rightarrow}{p}(\boldsymbol{x}(t))$ is the analytical score of the forward process that we described in Section˜D.3. In addition to $\boldsymbol{x}(t)$, we feed as input to $u_{\theta}$ a sinusoidal time embedding of dimension $32$. By initializing the weights of the output layer of $u_{\theta}$ to zero, we ensure that $p_{\theta}$ is initialized at $p_{\theta_{\text{base}}}$, the reversal of the forward diffusion process $\overset{\rightarrow}{p}$. We perform sampling using $10^{2}+1$ timesteps on a uniform time grid.
@@
Hyperparameter Genie2 ESM3-open TarFlow Gemma-2-9B-IT
@@
$\boldsymbol{x}$ Space $(\mathbb{R}^{100\times 3})^{100}$ $(\mathcal{S}^{100}_{4096})^{50}$ $\mathbb{R}^{256\times 256\times 3}$ $\mathcal{S}^{200}_{256{,}000}$
@@
Model parameters 1515M 1.4.4B 463463M 99B (110M trainable)
Training time (hrs) $48$ $2.3$ 33 66
E.2 Calibrating Genie2
@@
Since Genie2 is trained as the reversal of a discrete-time noising process (a DDPM, see Ho et al., 2020), we first convert the discrete-time denoising diffusion model to a (continuous-time) diffusion model. We achieve this by redefining the final timestep $T$ of the original denoising process to be time $1$ of the continuous-time process. To define the drift function, we take the DDPM transition mean defined at each time $t$ in the discrete-time process, divide it by $1/T=T$, and define the drift function to be equal to the resulting value in between times $t/T$ and $(t+1)/T$. The diffusion coefficient is similarly defined by the DDPM transition standard deviation at each time $t$ in the discrete-time process, but is instead scaled by $T^{1/2}$. This approach of converting the DDPM into a continuous-time diffusion model ensures that when the SDE is solved under the Euler-Maruyama scheme using a grid of $T$ timesteps (i.e., the original time grid used to define the DDPM), one samples from the original DDPM.
@@
For CGM-relax, we calibrate to k=99k{=}99 constraints on the bivariate CDF of alpha helix and beta strand content. And for CGM-reward, we calibrate to k=15k{=}15 constraints using $N=2.5\times 10^{4}$ samples from $p_{\theta_{\text{base}}}$. Since sampling from the Genie2 base model is time intensive, the sampling cost to compute $\widehat{\boldsymbol{\alpha}}_{N}$ (with small variance) is a downside to CGM-reward. Results reported in Figure˜4 are averages over $3$ trials, with two standard errors. We report results for CGM-relax with additional values of $\lambda$ in Figure˜7.
@@
Secondary structure annotation. As discussed in Section˜4.1, we measure the diversity of a collection of protein structures by computing the proportion of residues that lie in each of the three protein secondary structure types. For the CATH domains and Genie2, we perform annotations using the Biotite package (Kunzmann and Hamacher, 2018), which considers only $C_{\alpha}$ backbone atoms. For the CATH proteins (Sillitoe et al., 2021), we obtain the secondary structure distribution by annotating the domains collected and published by Ingraham et al. (2019).
Figure 7: Results from varying the regularization hyperparameter $\lambda$ in CGM-relax for Genie2 and ESM3. We observe that CGM-relax is insensitive to the choice of $\lambda$, with the exception of $\lambda=10^{1}$ for ESM3, in which case the fine-tuned model remains close to the base model.
@@ E.4 Calibrating TarFlow
Our motivation for choosing this problem was twofold. First, the quality of images generated by the base TarFlow model is high, such that a pre-trained classifier could attain high accuracy without fine-tuning. Second, we observe that the wildlife images generated by the base TarFlow model contained predominantly lions and leopards (Figure˜10), and rarely contained foxes or wolves. From $5\times 10^{3}$ samples annotated by GPT o5-mini, we computed animal proportions $\{0.3590,0.1260,0.0404,0.0256,0.2704,0.1752\}$.
@@
Although we require the model to state its confidence when labeling the images, we do not use these confidence scores for fine-tuning. The calibration function $\boldsymbol{h}(\boldsymbol{x})\in\mathbb{R}^{5}$ is a one-hot encoding of the first five classes. Since nearly all of the samples from the base TarFlow model are labeled as one of the six animals, we observe that choosing a six-dimensional constraint (i.e., adding cheetah) results in a poorly conditioned dual problem (7), since then the components of $\boldsymbol{h}$ are nearly linearly dependent. Our target is the uniform distribution over animals $\boldsymbol{h}^{\ast}=[.167,.167,.167,.167,.167]$.
@@
We perform calibration for $\lambda\in\{10^{-2},10^{-3},10^{-4},10^{-5}\}$, and sample size $N=5\times 10^{3}$. We report results as the mean over three replicates, with two standard errors. We assess the success of calibration using two metrics: the total variation (TV) distance of the distribution over animal proportions (using $5\times 10^{3}$ annotated images) to the uniform distribution; and the FID, computed using only the samples belonging to the wildlife class in the AFHQ training dataset. We use $5\times 10^{4}$ samples from the generative model to compute FID. It is important to note that the FID is an imperfect metric for assessing the quality of generated images since it will be lower for models whose animal class makeup is similar to that of the training distribution. To account for this, we evaluate CGM-relax on the maximum entropy reweighting of the training dataset to the uniform distribution over animal classes. In other words, we up or down weighted images belonging to a particular animal class in order to sample the six animals belonging to wildlife class with equal probability.
@@
Our best model, calibrated using CGM-relax with $\lambda=10^{-4}$, obtains class proportions $\{0.2248,0.0854,0.1750,0.1566,0.1668,0.1086\}$, evaluated using $5\times 10^{3}$ samples from the model; $0.0828$ of the samples were labeled as None. CGM-relax reduces the miscalibration error by three times, from a TV distance of .306.306 to .101.101 (Figure˜11). However, the FID score increases from $15.9$ to $21.0$. CGM-reward is unsuccessful at calibrating the base model; both miscalibration and FID is roughly the same as the base model. Since CGM-reward remains close to the base model, we evaluate FID on the original AFHQ training dataset.
@@
In Figure˜10, we provide random generations from both the pre-trained and the model calibrated with CGM-relax with $\lambda=10^{-4}$. By examining samples from the calibrated model, we observe two axes along which sample quality worsens after calibration. First, some of the samples (those labeled as None) are dogs or cats, which lie outside the AFHQ wildlife class. Second, a greater proportion of samples depict blends of multiple animals.
@@
Figure 11: Calibrating TarFlow with CGM-relax reduces the TV distance of animal class labels to the uniform distribution by three times. However, CGM-relax also produces fewer realistic samples, as measured by FID.
E.5 Calibrating Gemma-2-9B-IT
E.5.1 Autoregressive sampling and log-likelihoods
@@
We sample in the standard autoregressive fashion with no temperature scaling. To compute sequence log-likelihoods, we consider the prompt as given and ignore tokens generated after the first end-of-sequence (EOS) token. Let $m$ be the length of the prompt and $n$ the index of the first EOS token. Then sequence $\boldsymbol{x}$ has log-probability
@@
E.5.2 Prompt definition
We used the prompt “Write a short story (3-5 paragraphs) which only uses very simple words that a 3 year old child would likely understand. ONLY write the story without any additional text. Try to use characters with roughly EQUAL PROBABILITIES male or female. The story begins: "Once upon a time there was a <profession> named"”, which is similar to the one used by Eldan and Li (2023) with the addition of the EQUAL PROBABILITIES sentence. As shown in Figure˜12, we found that this addition decreased but did not eliminate gender imbalance compared to the original prompt.
Figure 12: Gender imbalance for profession stories generated by Gemma-2-9B-IT with and without gender balance prompt. Values are computed using $2,048$ samples per profession. Error bars are two times standard error of the mean approximated using the binomial variance.
E.5.3 Gemma-2-9B-IT constraint definition
To calibrate Gemma-2-9B-IT, we use a simple heuristic procedure to detect the gender of the story’s character associated with the profession in the prompt. The procedure returns $1$ for female, 1-1 for male, and 0 if the gender cannot be determined. Given a generated story, our procedure is as follows.
(1) Pronoun at sentence two. If the second sentence begins with a third-person singular pronoun, we assign gender based on that pronoun. This is common with our prompt template, e.g., “Once upon a time there was a doctor named Sam. She was very kind…”.
@@ E.5.3 Gemma-2-9B-IT constraint definition
E.6 LoRA details
We used LoRA with rank 32, scaling factor $\alpha=64$ and no LoRA dropout. LoRA was applied to the self-attention key, query, value, and output layers and to the MLP (gate, up, down) projection layers, resulting in $1.1\times 10^{8}$ trainable parameters.
E.6.1 Gemma-2-9B-IT figure details
Both panels of Figure˜6 were created using five replicates per model with different Pytorch seeds, with points indicating the mean metric value across replicates. $2048$ stories were sampled per profession for each replicate, resulting in $14\times 2048=28672$ stories per model. Due to low variance, two-times-standard-error-of-the-mean error bars are smaller than most markers, so are not shown.
@@ E.6.1 Gemma-2-9B-IT figure details
which takes maximum value $1$ if all samples used the same gender and minimum value 0 if there are an equal number of each gender. The overall miscalibration values shown on the y-axis of Figure˜6A were computed by taking the average miscalibration for the eight professions used for fine-tuning.
@@
This procedure was repeated five times for different sampling seeds to ensure variance from estimating $\boldsymbol{\alpha}$ had little effect on the outcome. The resulting error bars for the symmetrized KL are smaller than the marker size.
E.6.2 Example generations
@@
We provide example generations for four prompts before and after fine-tuning with CGM-relax with $\lambda=0.1$.
Listing 1: Samples from Gemma-2-9B-IT
Once upon a time there was a doctor named Ben. Ben had a big car. Ben had a red hat. He drove to the park. The park had a swing. A girl was on the swing. The girl had a red ball. The girl threw the ball. The ball went far. Ben smiled. Ben waved. The girl waved back. Ben went home.
Once upon a time there was a lawyer named Ben. Ben liked to wear a big tie. Ben went to work in a tall building. He saw lots of people in the big building. Ben sat at a big desk. He read papers all day. One day a girl named Sue came to see Ben. Sue had a puppy. Sue's puppy was lost. Sue was sad. Ben helped Sue. He said, "Don't worry, Sue. We will find your puppy." Ben looked all over the big building. He asked all the people. Finally, Ben found the puppy under a chair. The puppy was sleepy. Ben gave the puppy back to Sue. Sue was happy. Ben was happy.
Once upon a time there was a teacher named Mary. Mary had many students. They liked to play games. One day, Mary had a new game. It was a red ball. Mary said, "Roll the ball!" The students rolled the ball. It was fun! They rolled the ball to each other. They laughed so much. They played with the ball all day.At the end of the day, Mary said, "Time to clean up." The students helped Mary put the ball away. They said, "Thank you, Mary!" They liked playing games with her. Mary waved goodbye. She saw the students go home. Mary was happy. She liked teaching the students.
Once upon a time there was a pilot named Ben. He had a big red plane. Ben loved to fly. He would fly high in the sky. He saw birds and clouds. Once, Ben saw a big, shiny star. It was yellow and so bright. Ben smiled. He liked his plane. He liked to fly. One day, Ben went to the park. He saw a little girl. Her name was Lily. Lily looked at the sky. "I want to fly!" she said. Ben smiled. "You can fly in my plane!" he said. Lily was so happy. She got in the plane. Ben took off. Lily saw the clouds up close. They were white and fluffy. She laughed. Ben flew Lily high, high in the sky. Lily saw trees and cars. They looked small from up high. Ben and Lily flew for a long time. Then, Ben landed the plane. Lily hopped out. "Thank you!" she
@@
nce upon a time there was a doctor named Sam. Sam was nice. Sam had a toy red car. Sam liked to play with the car. He played with it all day. One day, Ben knocked on Sam's door. "Come in!" said Sam. Ben was sad. "Zoom zoom!" Ben saw the car. Ben liked cars too. "Can I play?" asked Ben. Sam said yes. They played with the car. They went zoom zoom all over the room. They played until it was time for Ben to go home. Ben said "Bye Sam! Thank you!" Ben ran home. Sam was happy. Sam liked to make friends happy.
Once upon a time there was a lawyer named Sam. Sam had a big, red car. He liked to drive fast. He went to work in his car every day. At work, Sam helped people. He talked to them. He wrote things down. He made plans. People liked Sam. He was nice. He made them feel better. Sam liked his job. One day, Sam saw a dog on the street. The dog was small and sad. Sam stopped his car. He got out and looked at the dog. Sam gave the dog a cookie.The dog wagged its tail. Sam smiled. He took the dog home with him. He named the dog Spot. Sam and Spot were happy.
Once upon a time there was a teacher named Ray. Ray had lots of toys. Ray liked to play with the toys.One day, a girl named Lily came to Ray's room. Lily wanted to play, too. Ray said, "Come play!" Ray got a big red ball. Lily liked the red ball. They played with the ball all day. They rolled it and threw it. They had lots of fun. After they played, Lily said, "Thank you, Ray!" Ray said, "You're welcome, Lily!" Lily went home. Ray put the toys back. Then Ray sat down and read a book. The end.
Once upon a time there was a pilot named Mia. Mia flew a big plane. She flew high in the sky. Mia saw a bird. The bird was small. Mia waved to the bird. The bird flew away. Mia saw a cloud. The cloud was white. Mia touched the cloud. The cloud was soft. Mia flew back home. Mia's dog was happy. Mia gave her dog a hug. Mia's cat was sad. Mia gave her cat a pet. " Mia liked to fly. She liked her dog and cat too.